1. The following is the demand function for a flavored milk producing company.


Q  =  5000 – 100 P


Where Q is the sales and P is the price of the drink.


a.     How many drink bottles could be sold at Rs.15 each?

b.    What should the price be for the company to sell 2000 drink bottles.


1
Expert's answer
2021-02-09T07:26:39-0500

ANSWER

a) When the price of the drink is equal to Rs.15 per bottle, customers can buy the 3,500 bottles according to the demand function. The following calculations were made:

"Q(15)=5,000-100*15=5,000-1,500=3,500"

b) The price that allows to sell 2,000 bottles can be calculated by means of demand function. Let's rewrite the demand function as follows:

"P(Q)=50-Q\/100"

Hence, "P(2,000)=50-2,000\/100=30"